# MasconCube: Fast and Accurate Gravity Modeling with an Explicit Representation

### v2

@@ 4.2 Hollow Body Modeling Challenges
Hollow body gravity inversion is inherently difficult due to Newton’s Shell theorem and the ill-posed nature of the inverse problem. According to Newton’s Shell theorem, a spherically symmetric hollow shell produces the same external gravitational field as a point mass at its center, making internal cavities virtually undetectable from external observations alone [31]. This fundamental ambiguity is exacerbated for concentric configurations, where multiple density distributions can produce nearly identical gravitational signatures. The Planetesimal - Hollow configuration features a central spherical cavity. As shown in Figure LABEL:fig:density_hollow, MasconCube does not outperform GeodesyNet, with both methods struggling to reconstruct the internal void.
We also tested an asymmetrical variant with a displaced hollow region, representing a theoretically less ambiguous case. As shown in Figure LABEL:fig:density_hollow_dec, while both methods successfully detect the asymmetry, neither adequately reconstructs the hollow region itself. MasconCube shows particularly limited capability in resolving internal void structures, suggesting that explicit mascon representations require additional constraints for complex internal geometries.
These results demonstrate a fundamental limitation in gravity-based density reconstruction: external gravitational observations effectively constrain surface density variations but provide insufficient information about internal cavities, especially symmetric ones.
@@ A.1 Tetrahedral Mesh Models
Table 7: Mass functions $m(x,y,z,V)$ used to generate the models, where a mascon is placed in correspondence to the centroid, with coordinates $x,y,z$ of each tetrahedral element with volume $V$. Everything is then normalized so that the total mass of the body sums to 1.
@@ A.2 Training Parameters
Parameter Value Notes
Epochs $1000$
@@
Learning rate $1\text{\times}{10}^{-5}$
Optimizer Adam $\beta_{1}=0.9,\beta_{2}=0.999$
Scheduling Step decay $\text{factor}=0.8\text{, patience}=200$
Table 8: Training hyperparameters used for MasconCube optimization across all experimental configurations.
@@ A.3 Trajectory Integration Parameters
Table 9: Physical parameters used in spacecraft trajectory simulations around each asteroid. Values for Itokawa, Bennu, and Eros are derived from real observational data, while planetesimal parameters are theoretical values based on plausible characteristics from planet formation theory. These parameters are essential for accurate modeling of gravitational forces, solar radiation pressure effects, and rotational dynamics in the trajectory integration process.

### v3

@@ 4.2 Hollow Body Modeling Challenges
Hollow body gravity inversion is inherently difficult due to Newton’s Shell theorem and the ill-posed nature of the inverse problem. According to Newton’s Shell theorem, a spherically symmetric hollow shell produces the same external gravitational field as a point mass at its center, making internal cavities virtually undetectable from external observations alone [26]. This fundamental ambiguity is exacerbated for concentric configurations, where multiple density distributions can produce nearly identical gravitational signatures. The Planetesimal - Hollow configuration features a central spherical cavity. As shown in Figure 21(c), MasconCube does not outperform GeodesyNet, with both methods struggling to reconstruct the internal void.
(c) Density reconstruction cross-sections for Planetesimal – Hollow showing central cavity. Both methods fail to accurately reconstruct the internal void structure.
We also tested an asymmetrical variant with a displaced hollow region, representing a theoretically less ambiguous case. As shown in Figure 21(d), while both methods successfully detect the asymmetry, neither adequately reconstructs the hollow region itself. MasconCube shows particularly limited capability in resolving internal void structures, suggesting that explicit mascon representations require additional constraints for complex internal geometries.
These results demonstrate a fundamental limitation in gravity-based density reconstruction: external gravitational observations effectively constrain surface density variations but provide insufficient information about internal cavities, especially symmetric ones. This limitation also has broader implications for the study of asteroid porosity. Since macroporosity and distributed voids influence the bulk density in ways that can be gravitationally degenerate with uniform-density models, accurate porosity estimation requires integrating gravitational data with complementary observations. Our findings thus emphasize that purely gravity-based inversions can reveal surface and bulk mass anomalies but remain insensitive to internal porosity distributions, an important consideration for planetary modeling and mission data interpretation [2, 5, 27].
(d) Density reconstruction cross-sections for Planetesimal – Hollow Decentered with off-center cavity. Methods detect asymmetry but struggle with void reconstruction.
@@ A.1 Tetrahedral Mesh Models
Table 8: Mass functions $m(x,y,z,V)$ used to generate the models. Each mascon is placed at the centroid $(x,y,z)$ of a tetrahedral element with volume $V$, and masses are normalized so the total body mass equals 1.
@@ A.2 Training Parameters
Table 9: Complete hyperparameter settings for MasconCube training.
Hyperparameter Value
Grid Configuration
Grid resolution $100\times 100\times 100$
Number of active mascons ($\hat{N}$) $10^{6}$ (body-dependent)
Initialization
[…]
Position initialization Fixed grid points
Random seed 40
Training Configuration
Total training steps 1000
@@
Batch refresh frequency Every 10 epochs
Sampling region Sphere of radius 1.0
Optimizer Settings
Optimizer Adam [17]
Initial learning rate $1\times 10^{-5}$
Learning rate schedule Step decay
Step decay factor 0.8
Step decay intervals 200
$\beta_{1}$ (Adam) 0.9 (default)
$\beta_{2}$ (Adam) 0.999 (default)
$\epsilon$ (Adam) $1\times 10^{-8}$ (default)
Weight decay 0
Loss Function
Loss type Normalized L1 (Equation 2)
Scaling factor computation Closed-form optimal $c$
Regularization
Mass normalization constraint $\sum_{j}m_{j}=1$ (enforced)
@@ A.3 Trajectory Integration Parameters
Table 10: Physical parameters used in spacecraft trajectory simulations. Values for Itokawa, Bennu, and Eros are from observations, while planetesimal parameters are theoretical, derived from planet formation models.